BexleyBrooks:

Need some more help on Math. 1. 4a-2a(a+9)=6 what is A? NO ARGUING OR SPAMMING I AM DONE WITH THAT!

HotPockets14:

Multiply the numbers\(\textcolor{#C58AF9}{1}\cdot \textcolor{#C58AF9}{4}a-2a(a+9)=6\)\(\textcolor{#C58AF9}{4}a-2a(a+9)=6\)2Distribute\(4a\textcolor{#C58AF9}{-2a(a+9)}=6\)\(4a\textcolor{#C58AF9}{-2a^{2}-18a}=6\)3Combine like terms\(\textcolor{#C58AF9}{4a}-2a^{2}\textcolor{#C58AF9}{-18a}=6\)\(\textcolor{#C58AF9}{-14a}-2a^{2}=6\)4Rearrange terms\(\textcolor{#C58AF9}{-14a-2a^{2}}=6\)\(\textcolor{#C58AF9}{-2a^{2}-14a}=6\)5Move terms to the left side\(-2a^{2}-14a=6\)\(-2a^{2}-14a-6=0\)6Common factor\(-2a^{2}-14a-6=0\)\(-2\left( a^{2}+7a+3\right) =0\)7Divide both sides by the same factor\(-2\left( a^{2}+7a+3\right) =0\)\(a^{2}+7a+3=0\)8Use the quadratic formula\(x=\frac{-\textcolor{#F28B82}{b}\pm \sqrt{\textcolor{#F28B82}{b}^{2}-4\textcolor{#C58AF9}{a}\textcolor{#8AB4F8}{c}}}{2\textcolor{#C58AF9}{a}}\)Once in standard form, identify a, b, and c from the original equation and plug them into the quadratic formula.\(a^{2}+7a+3=0\)\(a=\textcolor{#C58AF9}{1}\)\(b=\textcolor{#F28B82}{7}\)\(c=\textcolor{#8AB4F8}{3}\)\(a=\frac{-\textcolor{#F28B82}{7}\pm \sqrt{\textcolor{#F28B82}{7}^{2}-4\cdot \textcolor{#C58AF9}{1}\cdot \textcolor{#8AB4F8}{3}}}{2\cdot \textcolor{#C58AF9}{1}}\)9SimplifyEvaluate the exponent\(a=\frac{-7\pm \sqrt{\textcolor{#C58AF9}{7^{2}}-4\cdot 1\cdot 3}}{2\cdot 1}\)\(a=\frac{-7\pm \sqrt{\textcolor{#C58AF9}{49}-4\cdot 1\cdot 3}}{2\cdot 1}\)Multiply the numbers\(a=\frac{-7\pm \sqrt{49\textcolor{#C58AF9}{-4}\cdot \textcolor{#C58AF9}{1}\cdot \textcolor{#C58AF9}{3}}}{2\cdot 1}\)\(a=\frac{-7\pm \sqrt{49\textcolor{#C58AF9}{-12}}}{2\cdot 1}\)Subtract the numbers\(a=\frac{-7\pm \sqrt{\textcolor{#C58AF9}{49-12}}}{2\cdot 1}\)\(a=\frac{-7\pm \sqrt{\textcolor{#C58AF9}{37}}}{2\cdot 1}\)Multiply the numbers\(a=\frac{-7\pm \sqrt{37}}{\textcolor{#C58AF9}{2}\cdot \textcolor{#C58AF9}{1}}\)\(a=\frac{-7\pm \sqrt{37}}{\textcolor{#C58AF9}{2}}\)\(a=\frac{-7\pm \sqrt{37}}{2}\)10Separate the equations To solve for the unknown variable, separate into two equations: one with a plus and the other with a minus.\(a=\frac{-7+\sqrt{37}}{2}\\ a=\frac{-7-\sqrt{37}}{2}\)11SolveRearrange and isolate the variable to find each solution\(a=\frac{\sqrt{37}-7}{2}\\ a=\frac{-\sqrt{37}-7}{2}\)8 more stepsShow less Solution \(a=\frac{\sqrt{37}-7}{2}\\ a=\frac{-\sqrt{37}-7}{2}\)

TheRizzard:

rizzard